# Key Words.: binaries: spectroscopic – binaries: close – stars: early-type – stars: massive – Magellanic Clouds1 Introduction

### v1

@@ 3 Radial velocities
To measure radial velocities (RVs) we have used the same tools as for the B-type stars in 30 Doradus (30 Dor, Villaseñor et al., 2021), which are now part of the MINATO (Massive bINaries Analysis TOols, Villaseñor et al., 2023) python repository222https://github.com/jvillasr/MINATO, in a module called ravel. Briefly, it is a spectral line profile fitting tool, which uses Gaussian or Lorentzian (in the case of Balmer lines) profiles to fit a set of spectral lines. In the early B-type stars’ spectra, there is a rich set of spectral lines covered by the wavelength range of FLAMES; He i 4026, 4144, 4388, 4471, and the Balmer lines H and H. There are also a few metal lines that are usually used in the analysis of later-type stars such as Si ii 4128-31 and Mg ii 4481. However, these are very weak in the stars in our sample due to their spectral range and due to the low metallicity of the SMC, thus we have not used them in our RV measurements. For the earliest types, He ii 4542 is commonly present, but it is weaker than the He i lines and therefore provides less accurate RVs. This is similar for the case of Si iii 4553, which is usually too shallow for precise RV measurements as a result of the SMC’s low metallicity. All mentioned lines are fitted, but ravel selects the ones used to compute a weighted mean RV for each epoch based on the errors of the Gaussian or Lorentzian fit using a median-absolute-deviation (MAD; Hampel, 1974) test. For this reason, in most cases, only the He I and Balmer lines (H, H) are used, and Si II and Mg ii lines did not pass the MAD test for any of the systems. An example fit for a typical SB1 system is shown in the top plot of Fig. 3. Two He I lines and one Balmer line are displayed to illustrate the use of Gaussian and Lorentzian profiles.
@@ 3.1 Double-lined spectroscopic binaries
These enhancements lead to more reliable identification of SB2 components, improvements in RV uncertainties determination, and reduced the need of rejecting epochs due to misidentifications. However, challenges remain, particularly in cases with limited epochs, contributions from third objects, or significant profile variability. Future work will aim to refine our models to better address these complexities. The bottom plot of Fig. 3 demonstrates the SB2 fitting method by showcasing a double-component fit for the He i 4144 and He i 4388 lines of star BLOeM 1-055. The central panel (epoch 6) displays a phase near conjunction, where the two components are fully blended. In contrast, epochs with higher RV variations (e.g. epochs 5 and 7) provide better constraints on the line profiles of the individual components, enabling a reliable fit even for highly blended epochs.
@@ 3.2 Binary criteria
After weak or absent lines are omitted and epochs with a very low S/N () are rejected (see details in Villaseñor et al., 2021), we measure the peak-to-peak radial velocity difference (RV), which is the maximum difference between any two of the available nine epochs. These measurements are used to classify stars as spectroscopic binaries based on the criteria used by Sana et al. (2013); Dunstall et al. (2015); Bodensteiner et al. (2021); Banyard et al. (2022). The first criterion is based on the peak-to-peak RV variability. Stars with RV are classified as binary candidates, where is a specified threshold. Dunstall et al. (2015) adopted a value of km s-1 for the B-type stars in the LMC, based on their analysis of the binary fraction versus RV threshold, which showed a distinct break at 16 km s-1 for the B supergiants in their sample. Given that this feature is absent from our results, we adopt a value of km s-1, consistent with other studies (Sana et al., 2013; Bodensteiner et al., 2021; Banyard et al., 2022).
Furthermore, it is well known that B-type stars often exhibit pulsations (Aerts et al., 2009; Bowman, 2020). These pulsations can cause intrinsic RV variability with amplitudes peaking in the 10–20 km s-1 range (Stankov & Handler, 2005; Hey & Aerts, 2024). By choosing km s-1, we aim to exclude most stars displaying intrinsic variability caused by pulsations, thereby reducing contamination in our binary sample.
@@
where and are the individual RV measurements taken at two different epochs, and and are the uncertainties associated with each RV measurement. We define , representing the detection significance, which quantifies how significantly different two RV measurements are relative to their combined uncertainties. Since the RV uncertainties are critical for the determination of , we test their reliability in Appendix B. If either of the two criteria are not met, we classify the star as a radial velocity variable (RV var) if , otherwise, we do not consider the variability significant and classify those stars as radial velocity constant (RV cst). Stars classified as RV var inevitably include unidentified binaries in long-period orbits or with low-mass companions, in addition to stars showing intrinsic RV variability (such as pulsations or atmospheric variability), which may or may not be truly single stars. Due to this reason we correct for the observational biases and compute the intrinsic binary fraction (see Sect. 4).
For the binary classification procedure, we have first identified SB2 systems by visual inspection, and flagged them as SB2s so that ravel computes their RVs adequately. With the RV measurements computed, the binary classification is done automatically according to the two criteria, but some classifications were reevaluated by visual inspection of the fits to spectral line profiles and refitted. In the same way, we have classified three systems as SB3: BLOeM 5-062, BLOeM 6-062, and BLOeM 7-057. We base these classifications on the presence of 1 or 2 narrow cores in the He i lines, with broad profiles on both sides of the narrow lines, or a clear SB2 that cannot be successfully fitted with 2 Gaussian profiles (see examples in Appendix C.1). Furthermore, we found potential signals of contribution from a third star in the spectra of 15 SB2 systems. These have been flagged in the comments of Table LABEL:tab:bin (see also notes in Appendix C.2)
@@ 4 Multiplicity fraction
Given that our sample includes giant stars (luminosity class III) it is certainly possible that interactions might have already occurred. Furthermore, a lower surface gravity value, or analogously a larger radius, will affect the minimum orbital period before interaction () with consequences for the distribution of and the multiplicity fraction (Badenes et al., 2018). To quantify the effect of including moderately evolved stars, we compute (following Mazzola-Daher et al., 2022) for a range of primary masses (), surface gravities (), and mass ratios (), representative of our sample of B-type stars. From , we compute the allowed values for a range of orbital inclinations between 20 and 90 degrees. Figure 5 (top panel) shows how increases significantly towards lower values for a fixed mass ratio of , illustrating that systems with orbital periods below 5 d would have interacted once the primary reaches . Using would shift the critical periods roughly by -0.2 dex in . In the bottom panel we show the corresponding maximum peak-to-peak RV amplitude ( ) for a typical inclination of . Even at moderate inclinations and , the minimum remains around 100 km s-1 for most of our parameter space, safely above the 20 km s-1 threshold used in our binary criterion. Only in extreme cases of very low mass ratios () and low orbital inclinations () would fall below 20 km s-1 (white dashed contours). Therefore, even though the larger radii of giant stars shift the shortest detached periods to higher values, in most cases the resulting RV amplitudes remain well above our detection threshold.
@@
Here we defined the multiplicity fraction as that of the B-type primaries in our sample. Due to the unknown nature of most of the companions, we do not include them in our binary statistics. The multiplicity fraction quoted henceforth is effectively a close binary fraction that takes into account binaries at separations au, or , which roughly corresponds to the separations that will allow for binary interactions to happen. Consequently, this is the range of orbital periods for which observational biases are accounted for in the determination of the intrinsic bianry fraction (see Sect. 4.2). The 153 binaries in our sample yield an observed fraction of % , where the uncertainty represents the standard error of a proportion based on binomial statistics, corresponding to a 68% confidence interval. This is a significantly larger observed fraction than that found for the 30 Dor “unevolved” stars (; Dunstall et al., 2015). However, the detection probabilities of both surveys need to be carefully taken into account for a meaningful comparison. For example, the nine epochs analysed in this work provide an advantage over the six LR02 VFTS observations considered by Dunstall et al.. On the other hand, VFTS was much more sensitive to moderate-to-long period binaries due to its longer observational baseline of 10-12 months (and up to 22 months for 31 targets; Dunstall et al., 2015). We further discuss the differences between the samples of B-type stars and the implications for bias correction in Sect. 4.2.
@@
It is, however, possible to compare the observed multiplicity of our sample with that of the BLOeM O-type stars. Sana et al. (2024) determined an observed binary fraction of % for the 139 O-type stars in the BLOeM sample, which is in agreement with our findings for the B-type stars. This is interesting because the B-type population may suffer from different biases compared to the O-type stars due to their fainter magnitudes and lower S/N spectra. These biases result in a detection probability that drops at shorter orbital periods and larger companions masses (see Sana et al., 2024, and Sect. 6.1). A clear example of this is 30 Dor, where the observed multiplicity fraction was found to be for the O-type sample (Sana et al., 2013) and for the B-type stars (Dunstall et al., 2015). Given their faster evolution and higher detectability as binaries, the O-type population may be more affected by binary interactions. Binary interactions can effectively reduce the number of observable binaries through mergers or envelope stripping. Envelope stripping through stable mass transfer can lead to large mass ratios, making it difficult to detect RV variations and the flux contribution from the companion in optical wavelengths. Two possible scenarios might explain the similar observed multiplicity fractions of O- and B-type BLOeM stars: (1) the initial multiplicity fraction of O-type stars is higher than that of B-type stars in the SMC but has decreased due to binary interactions; or (2) both populations have similar initial multiplicity fractions, but binary interactions are less frequent during the early phases of evolution at low metallicity. While the first option seems the most logical, there is also evidence that might support the latter possibility. As a result of the significantly reduced atmospheric opacity compared to stars in the LMC and MW, MS massive stars in the SMC are more compact, hotter, and more luminous (Maeder & Meynet, 2000; Götberg et al., 2017), potentially delaying Roche-lobe overflow (Smith, 2014). Based on evolutionary models, Klencki et al. (2020) found that at metallicities comparable to that of the SMC, the most common post-MS donors in massive binaries ( 16–50) are slowly expanding core-He burning (CHeB) stars. This means that interactions take place on average at later phases of evolution due to the smaller radii of low-metallicity massive stars. Even if this effect is not large enough to make a difference during the MS, it might play an important role for stars crossing the Hertzsprung gap, during which case-B mass transfer occurs. Although not the scope of this work, binary population synthesis modelling could help evaluate the significance of metallicity effects across different evolutionary phases and their impact on the multiplicity fraction of a population of massive stars, such as that of the BLOeM OB-type sample.
@@
In Fig. 7 we show the dependence of our observed multiplicity fraction on the adopted value of the cut-off. We do not observe a significant feature that might indicate a transition between variability caused by intrinsic features (e.g. winds, pulsations) and that caused by orbital motion. This is in contrast to what have been observed for the O-type stars and B-type supergiants in 30 Dor (Sana et al., 2013; Dunstall et al., 2015), where a distinct “kink” is visible at 16 km s-1. It is not clear if this is a consequence of the less active atmospheres of the B-type stars at low metallicity or what the role of pulsations might be. However, since the winds of massive stars are much reduced at SMC metallicity (Vink et al., 2001; Björklund et al., 2021), we do not expect winds to play a major role in the RV variability of the B III/V stars (e.g. Ramachandran et al., 2019). Similarly, if pulsations were to strongly contribute to the RV variability, we would expect a feature as a kink close to 20 km s-1 where the fraction of large RV amplitude variations drops significantly. We will come back to the role of pulsations in Sect 6.1, where we discuss the impact of adopting different RV threshold values.
4.1 Variations of the observed multiplicity fraction
We have computed the observed binary fraction for each of the eight FLAMES fields in the BLOeM survey, as shown in Table 2. We do not find a significant variation across fields, with most of them exhibiting multiplicity fractions close to 50% within . The exception is field 2, which exhibits a multiplicity fraction of %. Compared to the mean fraction of the other 7 fields (% ), this deviation appears statistically significant ( and ). However, field 2 has a relatively small sample size () compared to the other fields. The high uncertainty increases the likelihood that the observed difference is due to statistical fluctuations inherent in small sample sizes. We have not identified any environmental factors unique to field 2 that would result in a higher multiplicity fraction. Therefore, we interpret this result cautiously and attribute the higher fraction in field 2 to small number statistics rather than a genuine variation in the underlying stellar population.
We have also investigated potential variations in the multiplicity fraction with spectral type, as a proxy for temperature. In the histogram of Fig. 8 we show the number of stars per spectral type and luminosity class, with their corresponding oberved multiplicity fraction annotated. To minimise the impact of small number statistics, spectral types have been grouped as follows: the B0 bin includes spectral types from B0 to B0.5, B0.7 is grouped with B1, and B1.5 objects remain ungrouped, as in Fig. 2. The few B2.5 stars have been incorporated into the B2 bin. The purple unfilled histogram represents the total number of stars in each spectral type bin, while denotes the total multiplicity fraction within each bin, corresponding to the number of stars in the stacked histogram. Uncertainties in multiplicity fraction were computed using binomial statistics. Furthermore, the number of stars per luminosity class within each spectral type group is shown as stacked segments, each with its multiplicity fraction. It is not possible to draw strong conclusions from Fig. 8 due to the small number of stars with luminosity classes V and IV. For the multiplicity fraction of giants and total fraction, there are no significant trends with temperature, suggesting that within our spectral range, the multiplicity fraction remains constant.

### v4

@@ 3 Radial velocities
To measure RVs we have used the same tools as for the B-type stars in 30 Dor (Villaseñor et al., 2021), which are now part of the MINATO (Massive bINaries Analysis TOols, Villaseñor et al., 2023) python repository333https://github.com/jvillasr/MINATO, in a module called ravel. Briefly, it is a spectral line profile fitting tool, which uses Gaussian or Lorentzian (in the case of Balmer lines) profiles to fit a set of spectral lines. In the early B-type stars’ spectra, there is a rich set of spectral lines covered by the wavelength range of FLAMES; He i 4026, 4144, 4388, 4471, and the Balmer lines H and H. There are also a few metal lines that are usually used in the analysis of later-type stars such as Si ii 4128-31 and Mg ii 4481. However, these are very weak in the stars in our sample due to their spectral range and due to the low metallicity of the SMC, thus we have not included them in our RV measurements. For the earliest types, He ii 4542 is commonly present, but it is weaker than the He i lines and therefore provides less accurate RVs. This is similar for the case of Si iii 4553, which is usually too shallow for precise RV measurements as a result of the SMC’s low metallicity. All mentioned lines are fitted, but ravel selects the ones used to compute a weighted mean RV for each epoch based on the errors of the Gaussian or Lorentzian fit using a median-absolute-deviation (MAD; Hampel, 1974) test. For this reason, in most cases, only the He I and Balmer lines (H, H) are used, and Si II and Mg ii lines did not pass the MAD test for any of the systems. An example fit for a typical SB1 system is shown in the top plot of Fig. 2. Two He I lines and one Balmer line are displayed to illustrate the use of Gaussian and Lorentzian profiles.
@@ 3.1 Double-lined spectroscopic binaries
These enhancements lead to more reliable identification of SB2 components, improvements in the determination of RV uncertainties, and reduced the need to reject epochs due to misidentifications. However, challenges remain, particularly in cases with limited epochs, contributions from third objects, or significant profile variability. Future work will aim to refine our models to better address these complexities. The bottom plot of Fig. 2 demonstrates the SB2 fitting method by showcasing a double-component fit for the He i 4144 and He i 4388 lines of star BLOeM 1-055. The central panel (epoch 6) displays a phase near conjunction, where the two components are fully blended. In contrast, epochs with higher RV variations (e.g. epochs 5 and 7) provide better constraints on the line profiles of the individual components, enabling a reliable fit even for highly blended epochs.
@@ 3.2 Binary criteria
After weak or absent lines are omitted and epochs with a very low S/N () are rejected (see details in Villaseñor et al., 2021), we measure the peak-to-peak RV difference (RV), which is the maximum difference between any two of the available nine epochs. These measurements are used to classify stars as spectroscopic binaries based on the criteria used by Sana et al. (2013); Dunstall et al. (2015); Bodensteiner et al. (2021); Banyard et al. (2022). The first criterion is based on the peak-to-peak RV variability. Stars with RV are classified as binary candidates, where is a specified threshold. Dunstall et al. (2015) adopted a value of km s-1 for the B-type stars in the LMC, based on their analysis of the binary fraction versus RV threshold, which showed a distinct break at 16 km s-1 for the B supergiants in their sample. Given that this feature is absent from our results, we adopt a value of km s-1, consistent with other studies (Sana et al., 2013; Bodensteiner et al., 2021; Banyard et al., 2022).
Furthermore, it is well known that B-type stars often exhibit pulsations (Aerts et al., 2009; Bowman, 2020). These pulsations can cause intrinsic RV variability with amplitudes peaking in the 10–20 km s-1 range (Stankov & Handler, 2005; Hey & Aerts, 2024). By adopting km s-1, we aim to exclude most stars displaying intrinsic variability caused by pulsations, thereby reducing contamination in our binary classification.
@@
where and are the individual RV measurements taken at two different epochs, and and are the uncertainties associated with each RV measurement. We define , representing the detection significance, which quantifies how significantly different two RV measurements are relative to their combined uncertainties. Since the RV uncertainties are critical for the determination of , we test their reliability in Appendix B. If either of the two criteria are not met, we classify the star as a RV variable (RV var) if , otherwise, we do not consider the variability significant and classify those stars as RV constant (RV cst). Stars classified as RV var inevitably include unidentified binaries in long-period orbits or with low-mass companions, in addition to stars showing intrinsic RV variability (such as pulsations or atmospheric variability), which may or may not be truly single stars. For this reason, we correct for the observational biases and compute the intrinsic binary fraction (see Sect. 4).
For the binary classification procedure, we have first identified SB2 systems by visual inspection, and flagged them as SB2s so that ravel computes their RVs adequately. With the RV measurements computed, the binary classification is done automatically according to the two criteria, but some classifications were reevaluated by visual inspection of the fits to spectral line profiles and refitted. In the same way, we have classified three systems as SB3: BLOeM 5-062, BLOeM 6-062, and BLOeM 7-057. We base these classifications on the presence of one or two narrow cores in the He i lines, with broad profiles on both sides of the narrow lines, or a clear SB2 that cannot be successfully fitted with two Gaussian profiles (see examples in Appendix C.1). Furthermore, we found potential signals of contribution from a third star in the spectra of 15 SB2 systems. These have been flagged in the comments of Table D.1 (see also notes in Appendix C.2).
@@ 4 Multiplicity fraction
Given that our sample includes giant stars (luminosity class III) it is certainly possible that interactions might have already occurred. A lower surface gravity value, or analogously a larger radius, will affect the minimum orbital period before interaction () with consequences for the distribution of and the multiplicity fraction (Badenes et al., 2018). To quantify the effect of including moderately evolved stars, we compute (following Mazzola-Daher et al., 2022) for a range of primary masses (), surface gravities (), and mass ratios (), representative of our sample of B-type stars. From , we compute the allowed values for a range of orbital inclinations between 20 and 90 degrees. Figure 4 (top panel) shows how increases significantly towards lower values for a fixed mass ratio of , illustrating that systems with orbital periods below 5 d would have interacted once the primary reaches . Using would shift the critical periods roughly by -0.2 dex in . In the bottom panel we show the corresponding maximum peak-to-peak RV amplitude ( ) for a typical inclination of . Even at moderate inclinations and , the minimum remains around 100 km s-1 for most of our parameter space, safely above the 20 km s-1 threshold used in our binary criterion. Only in extreme cases of very low mass ratios () and low orbital inclinations () would fall below 20 km s-1 (dashed white contours). Therefore, even though the larger radii of giant stars shift the shortest detached periods to higher values, in most cases the resulting RV amplitudes remain well above our detection threshold.
4.1 Observed multiplicity fraction
@@
Here we defined the multiplicity fraction as that of the B-type primaries in our sample. Due to the unknown nature of most of the companions, we do not include them in our binary statistics. The multiplicity fraction quoted henceforth is effectively a close binary fraction that takes into account binaries at separations au, or , which roughly corresponds to the separations that will allow for binary interactions to happen. Consequently, this is the range of orbital periods for which observational biases are accounted for when determining the intrinsic binary fraction (see Sect. 4.3). The 153 binaries in our sample yield an observed fraction of % , where the uncertainty represents the standard error of a proportion based on binomial statistics, corresponding to a 68% confidence interval. This is a significantly larger observed fraction than that found for the 30 Dor “unevolved” stars (; Dunstall et al., 2015). However, the detection probabilities of both surveys need to be carefully taken into account for a meaningful comparison. For example, the nine epochs analysed in this work provide an advantage over the six LR02 VFTS observations considered by Dunstall et al.. On the other hand, VFTS was much more sensitive to moderate-to-long period binaries due to its longer observational baseline of 10-12 months (and up to 22 months for 31 targets; Dunstall et al., 2015). We further discuss the differences between the samples of B-type stars and the implications for bias correction in Sect. 4.3.
@@
It is, however, possible to compare the observed multiplicity of our sample with that of the BLOeM O-type stars. Sana et al. (2024) determined an observed binary fraction of % for the 139 O-type stars in the BLOeM sample, which is in agreement with our findings for the B-type stars. This is interesting because the B-type population may suffer from different biases compared to the O-type stars due to their fainter magnitudes and lower S/N spectra. These biases result in a detection probability that drops at shorter orbital periods and larger companions masses for B-type stars (see Sana et al., 2024, and Sect. 6.1). A clear example of this is 30 Dor, where the observed multiplicity fraction was found to be for the O-type sample (Sana et al., 2013) and for the B-type stars (Dunstall et al., 2015). Given their faster evolution and higher detectability as binaries, the O-type population may be more affected by binary interactions. Binary interactions can effectively reduce the number of observable binaries through mergers or envelope stripping. Envelope stripping through stable mass transfer can lead to large mass ratios, making it difficult to detect RV variations and the flux contribution from the companion in optical wavelengths. Two possible scenarios might explain the similar observed multiplicity fractions of O- and B-type BLOeM stars: (1) the initial multiplicity fraction of O-type stars is higher than that of B-type stars in the SMC but has decreased due to binary interactions; or (2) both populations have similar initial multiplicity fractions, but binary interactions are less frequent during the early phases of evolution at low metallicity. While the first option seems the most logical, there is also evidence that might support the latter possibility. As a result of the significantly reduced atmospheric opacity compared to stars in the LMC and MW, MS massive stars in the SMC are more compact, hotter, and more luminous (Maeder & Meynet, 2000; Götberg et al., 2017), potentially delaying Roche-lobe overflow (Smith, 2014). Based on evolutionary models, Klencki et al. (2020) found that at metallicities comparable to that of the SMC, the most common post-MS donors in massive binaries ( 16–50) are slowly expanding core-He burning (CHeB) stars. This means that interactions take place on average at later phases of evolution due to the smaller radii of low-metallicity massive stars. Even if this effect is not large enough to make a difference during the MS, it might play an important role for stars crossing the Hertzsprung gap, during which case-B mass transfer occurs. Although not the scope of this work, binary population synthesis modelling could help evaluate the significance of metallicity effects across different evolutionary phases and their impact on the multiplicity fraction of a population of massive stars, such as that of the BLOeM OB-type sample.
In Fig. 6 we show the dependence of our observed multiplicity fraction on the adopted value of the cut-off. We do not observe a significant feature that might indicate a transition between variability caused by intrinsic features (e.g. winds or pulsations) and that caused by orbital motion. This is in contrast to what has been observed for the O-type stars and B-type supergiants in 30 Dor (Sana et al., 2013; Dunstall et al., 2015), where a distinct “kink” is visible at 16 km s-1. It is not clear if this is a consequence of the less active atmospheres of the B-type stars at low metallicity or what the role of pulsations might be. However, since the winds of massive stars are much reduced at SMC metallicity (Vink et al., 2001; Björklund et al., 2021), we do not expect winds to play a major role in the RV variability of the B III/V stars (e.g. Ramachandran et al., 2019). Similarly, if pulsations were to strongly contribute to the RV variability, we would expect a feature as a kink close to 20 km s-1 where the fraction of large RV amplitude variations drops significantly. We will come back to the role of pulsations in Sect 6.1, where we discuss the impact of adopting different RV threshold values.
4.2 Variations in the observed multiplicity fraction
We have computed the observed binary fraction for each of the eight FLAMES fields in the BLOeM survey, as shown in Table 2. We do not find a significant variation across fields, with most of them exhibiting multiplicity fractions close to 50% within . The exception is Field 2, which exhibits a multiplicity fraction of %. Compared to the mean fraction of the other seven fields (% ), this deviation appears statistically significant ( and ). However, Field 2 has a relatively small sample size () compared to the other fields. The high uncertainty increases the likelihood that the observed difference is due to statistical fluctuations inherent in small sample sizes. We have not identified any environmental factors unique to Field 2 that would result in a higher multiplicity fraction. Therefore, we interpret this result cautiously and attribute the higher fraction in Field 2 to small number statistics rather than a genuine variation in the underlying stellar population.
@@ 4.2 Variations in the observed multiplicity fraction
We have also investigated potential variations in the multiplicity fraction with spectral type, as a proxy for mass. In the histogram of Fig. 7 we show the number of stars per spectral type and luminosity class, with their corresponding observed multiplicity fractions annotated. To minimise the impact of small number statistics, spectral types have been grouped as follows: the B0 bin includes spectral types from B0 to B0.5, B0.7 is grouped with B1, and B1.5 objects remain ungrouped, as in Fig. 1. The few B2.5 stars have been incorporated into the B2 bin. The purple unfilled histogram represents the total number of stars in each spectral type bin, while denotes the total multiplicity fraction within each bin, corresponding to the number of stars in the stacked histogram. Uncertainties in multiplicity fraction were computed using binomial statistics. Furthermore, the number of stars per luminosity class within each spectral type group is shown as stacked segments, each with its multiplicity fraction. It is not possible to draw strong conclusions from Fig. 7 due to the small number of stars with luminosity classes V and IV. For the multiplicity fraction of giants and total fraction, there are no significant trends with temperature, suggesting that within our spectral range, the multiplicity fraction remains constant.
4.3 Intrinsic multiplicity fraction